Double your money -Rule of 72.  Approximately how long will it take to double your money if you get an annual return of 4.4 %​, 6.1 %​, or 11.7 % on your​ investment?  
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut copy ▼ Format Painter E AutoSum ー E ゴWrap Text Sort &Find & Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼鼎 Cells Edting Number Alignment Clipboard CE444 CH CE CF CD CA св 432 433 434 435 436 437 438 439 440 441 442 443 RULE OF 72 NO OF YEARS TO DOUBLE THE MONEY = 72/ RATE = NO OF YEARS TO DOUBLE THE MONEY = 72/ 4.4 = 16.36 YEARS NO OF YEARS TO DOUBLE THE MONEY = 72/ 6.1 11.80 YEARS NO OF YEARS TO DOUBLE THE MONEY = 72/ 11.7 6.15 YEARS 445 446 447 448 449 450 /4 | CALCULATOR , LOAN OPTIONS BOX DILUTION DOLLAR COST AVGMORTGAGE EXPO Sheeti Sheet2 Shel CAP STRU VALUE 福 130% 06:51 31-01-2019 rences: x261